Linux HOWTO Index

by Matt Welsh, mdw@sunsite.unc.edu v2.0, 17 March 1994

This is an index to the Linux HOWTO documents.

Contents

1 What are Linux HOWTOs? 1

2 Where to get Linux HOWTOs 1

3 HOWTO Index 2

4 Wanted HOWTOs 3

5 Writing and submitting a HOWTO 4

6 Miscellaneous and Legalese 4

1 What are Linux HOWTOs?

Linux HOWTOs are short online documents which describe in detail a certain aspect of configuring or using the Linux system. For example, there is the Installation HOWTO, which gives instructions on installing Linux, and the Mail HOWTO, which describes how to set up and configure mail under Linux. Other examples include the NET-2-HOWTO (previously the NET-2-FAQ) and the Printing HOWTO.

Information in HOWTOs is generally more detailed and in-depth than what can be squeezed into the Linux FAQ. For this reason, the Linux FAQ is being rewritten. A large amount of the information contained therein will be relegated to various HOWTO documents. The FAQ will be a shorter list of frequently asked questions about Linux, covering small specific topics. Most of the “useful” information in the FAQ will now be covered in the HOWTOs.

HOWTOs are comprehensive docs—much like an FAQ but generally not in question-and-answer format. However, many HOWTOs contain an FAQ section at the end. For example, the NET-2-FAQ has been renamed to the NET-2-HOWTO, because it wasn’t in question-and-answer format. However, you will see the NET-2-HOWTO named as the NET-2-FAQ in many places. The two docs are one and the same.

2 Where to get Linux HOWTOs

HOWTOs can be retrieved via anonymous FTP from the following sites:

• sunsite.unc.edu:/pub/Linux/docs/HOWTO
• tsx-11.mit.edu:/pub/linux/docs/HOWTO

as well as the many mirror sites, which are listed in the Linux META-FAQ (see below).

The Index, printed below, lists the currently available HOWTOs.

HOWTOs are also posted regularly to the newsgroups comp.os.linux and comp.os.linux.announce. In addition, a number of the HOWTOs will be crossposted to news.answers. Therefore, you can find the Linux HOWTOs on the news.answers archive site rtfm.mit.edu.

3 HOWTO Index

The following Linux HOWTOs are currently available. The list is small, because this is a relatively new project.

• Linux INFO-SHEET, by Michael K. Johnson (johnsonm@sunsite.unc.edu). A general technical introduction to Linux, listing its features, supported hardware, and other general information.

• Linux META-FAQ, by Michael K. Johnson (johnsonm@sunsite.unc.edu). A list of sources of information about Linux, such as documentation, newsgroups, FTP sites, and so on.

• Installation HOWTO, by Matt Welsh (mdw@sunsite.unc.edu). How to obtain and install Linux. This is the first document which a Linux newcomer should read. It discusses how to obtain Linux, both from the Internet and via Snail-mail, and how to install it. Focuses primarily on the SLS distribution of Linux, which is the current de facto standard.

• Linux Distribution HOWTO, by Matt Welsh (mdw@sunsite.unc.edu). A list of many of the major Linux distributions available via anonymous FTP and via mail order on diskette, tape, or CD-ROM. Also includes other Linux-related goodies that you can only get via mail order.

• NET-2 HOWTO, by Terry Dawson (terryd@extro.ucc.su.oz.au). Describes how to install and configure the Linux NET-2 TCP/IP code, in kernels 0.99.pl10 and above. (This has nothing to do with the so-called BSD NET-2 distribution.) Discusses various aspects of TCP/IP networking under Linux, including the use of SLIP. The complete instruction manual for TCP/IP under Linux.

• Ethernet HOWTO, by Paul Gortmaker (gpg109@rsphysse.anu.edu.au). A HOWTO about the various Ethernet device drivers which are available for Linux. Should be read along with the NET-2 HOWTO by anyone using Ethernet for TCP/IP on Linux.

• Serial HOWTO, by Greg Hankins (gregh@cc.gatech.edu). Information on serial communications, devices, and software packages for Linux.

• Term HOWTO, by Bill Reynolds (bill@goshawk.lanl.gov). Information on using the term communications package with Linux.

• XFree86 HOWTO, by Helmut Geyer (geyer@polyhymnia.iwr.uni-heidelberg.de). This

HOWTO describes how to install and configure the X-Windows software distribution for Linux. Most of this is Linux-specific, and expects you to read the documentation included in the XFree86-2.0 package.

• MGR HOWTO, by Vincent Broman (broman@Np.nosc.mil). A HOWTO describing the MGR window system for Linux, another graphical interface somewhat like X.

• Mail HOWTO, by Vince Skahan (vince@victrola.wa.com). All about configuration of electronic mail with Linux, including UUCP and TCP/IP mailers such as Smail and sendmail.

• UUCP HOWTO, by Vince Skahan (vince@victrola.wa.com). How to configure UUCP to transfer files, mail, and news to other UNIX machines on a UUCP-based connection. Includes discussion of Taylor UUCP for Linux.

• News HOWTO, by Vince Skahan (vince@victrola.wa.com). Setup and configuration of news software, such as C News, INN, and newsreaders (such as rn and tin) under Linux.

• Printing HOWTO, by Grant Taylor (gtaylor@cs.tufts.edu). How to configure and use printing software under Linux, such as lp, lpr, etc. including remote printing.

• SCSI HOWTO, by Drew Eckhardt (drew@kinglear.cs.Colorado.EDU). All about SCSI tapes, drives, and CD-ROMs under Linux.

• Hardware Compatibility HOWTO, by Ed Carp (erc@apple.com). An extensive list of hardware supported by Linux.

• Ftape HOWTO, by Kai Harrekilde-Petersen (ftape@mic.dth.dk). A HOWTO describing the setup and use of various QIC ftape drivers for Linux.

• Sound HOWTO, by Jeff Tranter (tranter@software.mitel.com). Information on sound-related hardware and software for Linux, such as how to configure your system for use with a sound card.

• JE HOWTO, by Yasuhiro Yamazaki (hiro@rainbow.physics.utoronto.ca). Information on JE, a set of Japanese Extensions for Linux, including software to edit, display, and print Japanese text.

• Busmouse HOWTO, by Mike Battersby (mike@starbug.apana.org.au). Information on using busmice with Linux and XFree86.

• Tips HOWTO, by Vince Reed (reedv@rpi.edu). Miscellaneous tips and tricks for Linux, which don’t fit well into any of the other HOWTOs.

• Keystroke HOWTO, by Zenon Fortuna (zenon@netcom.com). Information on binding macros to keystrokes under Linux.

4 Wanted HOWTOs

I would be very happy if someone would write the following HOWTO documents, and submit them to me (see the instructions in the next section below). Most of these are modifications of existing FAQ documents, and should be easy to put into HOWTO format. Keep in mind that HOWTOs generally aren’t in FAQ format, but can contain an FAQ section at the end.

Some of these are under construction; however, if you’re interested in working on one of them, please let me know.

• GCC/LIBRARY HOWTO. Discussing installation/upgrading of the GCC and shared libraries under Linux. How to build shared libraries. Most of the material would come from Mitchum DeSouza’s GCC FAQ.

• KERNEL HOWTO. How to upgrade and compile the kernel. Basic kernel tweaking, including how to turn on debugging output. How to decrypt panic messages. Basic introduction to kernel structure and hacking.

• If you have an idea for another HOWTO, please mail me!

5 Writing and submitting a HOWTO

If you would like to write a Linux HOWTO document, there are a few guidelines that you should follow.

• Format the document neatly. HOWTOs must be available in plain ASCII format, but you are free to use a formatting tool (texinfo, LameTeX, nroff) to format the document. Try to use meaningful structure and organization, and write clearly. Remember that many of the people reading HOWTOs do not speak English as their first language.

• You may wish to use the Linuxdoc-SGML package, available from ftp.cs.cornell.edu:/pub/mdw/, to format the HOWTO. This package allows you to produce LaTeX, plain ASCII, and HTML from a single source document, and was designed specifically with the HOWTOs in mind. If you are writing a new HOWTO you should seriously consider this.

• Make sure that all of the information is correct. I can’t stress this enough. When in doubt, speculate, but make it clear that you’re only guessing.

• Make sure that you are covering the most recent version of the available software. Also, be sure to include full instructions on where software can be downloaded from (FTP site name, full pathname).

• Include an FAQ section at the end, if appropriate. Many HOWTO documents need an ”FAQ” or ”Common Problems” section to cover information which can’t be covered in the regular text.

• Use other HOWTOs as a model.

After you have written the HOWTO, mail it to me (Matt Welsh) at mdw@sunsite.unc.edu. If you have use Linuxdoc-SGML, simply mail me the SGML source; I take care of formatting the documents. I’ll also take care of archiving the HOWTO and posting it to the various newsgroups.

It is important that you go through me when submitting a HOWTO, as I maintain the archives and need to keep track of what HOWTOs are being written and who is doing what.

All you have to do is send me periodic updates (every month or so, as needed) and I will take care of posting the HOWTO regularly.

6 Miscellaneous and Legalese

If you have questions, please feel free to mail mdw@sunsite.unc.edu. The Linux FAQ rewrite is being coordinated by Ian Jackson, ijackson@nyx.cs.du.edu, with help from others.

Unless otherwise stated, Linux HOWTO documents are copyrighted by their respective authors. Linux HOWTO documents may be reproduced and distributed in whole or in part, in any medium physical or electronic, without permission of the author. Translations and derivative works are similarly permitted without express permission. Commercial redistribution is allowed and encouraged; however, the author would like to be notified of any such distributions.

In short, we wish to promote dissemination of this information through as many channels as possible. However, we do wish to retain copyright on the HOWTO documents, and would like to be notified of any plans to redistribute the HOWTOs. If you have questions, please contact Matt Welsh, the Linux HOWTO

coordinator, at mdw@sunsite.unc.edu.